Assume the following demand and supply equations:

Demand: Q = 480 - 35P

Supply: Q = 200 +16P

What is the equilibrium Price? What is the equilibrium quantity?

What will happen if price is fixed at 3?
Answer #1

Answer

the equilibrium is at Qd=Qs

480-35P=200+16P

51P=280

P=5.49019608

Q=480-35*5.49019608=287.843137

The price is $5.49 and the quantity is 287.84.

======

P=3 then

Qd=480-35*3=375

Qs=200+16*3=248

Qd>Qs so there is a shortage as the price is below equilibrium price so the quantity demanded is higher than the quantity supplied.

Shortage =Qd-Qs=375-248=127

The shortage is 127 units